OpenStudy (pottersheep):

is cos^3 + sin^3 = 1?

myininaya (myininaya):

\[\cos^2(\theta)+\sin^2(\theta)=1\]

OpenStudy (anonymous):

not usually

OpenStudy (pottersheep):

Okay thank you :)